OpenStudy (anonymous):

Write the equation of the line in slope-intercept form that passes through the point (3, -2) and has a slope of 2.

OpenStudy (aroub):

y=mx+b , m=2 , (3,-2) -2=2(3)+b -2=6+b -6-2=b -8=b y=2x-8